### v1

# On-Policy Optimization of ANFIS Policies Using Proximal Policy Optimization

Abstract
We propose a reinforcement learning (RL) approach for training neuro-fuzzy controllers using Proximal Policy Optimization (PPO). Building on prior work that applied Deep Q-Learning to Adaptive Neuro-Fuzzy Inference Systems (ANFIS), our method replaces the off-policy value-based framework with a stable on-policy actor–critic loop. We evaluate this approach in the CartPole-v1 environment using multiple random seeds and compare its learning performance against ANFIS-Deep Q-Network(DQN) baselines. It was found that PPO-trained fuzzy agents achieved a mean return of $500\pm 0$ on CartPole-v1 after 20000 updates, showcasing less variance than prior DQN-based methods during training and overall faster convergence. These findings suggest that PPO offers a promising pathway for training explainable neuro-fuzzy controllers in reinforcement learning tasks.
Keywords:
Fuzzy Logic Optimization PPO Explainable AI Trustworthy AIInterpretability
1 Introduction
Deep reinforcement learning (RL) has shown the potential to display super-human skill in complex domains. An example of this is AlphaGo’s defeat of a Go world champion [1]. However, the policies learned by deep neural networks (DNNs) remain largely opaque, limiting trust in safety-critical settings such as autonomous driving and healthcare.
In contrast, fuzzy inference systems offer transparency while providing a robust solution . These systems fall into two families: Mamdani and Takagi–Sugeno–Kang (TSK). Mamdani systems rely on linguistic IF–THEN rules with fuzzy outputs and subsequent defuzzification, making them highly interpretable but less amenable to gradient-based tuning [2]. TSK models instead express rule consequents as algebraic functions of the inputs, producing smoother numeric outputs and enabling more robust numerical optimization [3].
Yet both architectures still lack systematic training pipelines. Designing membership functions, rule bases, and consequents typically depends on expert heuristics or search methods such as genetic algorithms, which hampers scalability to high-dimensional or dynamic tasks [4].
Neuro-fuzzy methods like ANFIS address this by using a neural network to transform the inputs into intermediate features, which feed into Gaussian membership functions that activate first-order TSK rules; their weighted outputs are aggregated to produce the final action logits. All trainable parameters like network weights, membership centres and sigmas, and rule consequents are updated through gradient descent [5]. Furthermore, deep applications like ANFIS-DQN hybrids have shown promise [6] but inherit the instability of off-policy Q-learning.
Proximal Policy Optimization (PPO) combats these issues with a clipped, on-policy surrogate objective that yields stable learning and strong sample efficiency [7]. We therefore integrate an ANFIS-style fuzzy module into PPO, forming a PPO-Fuzzy agent. Using the well-studied CartPole-v1 benchmark, we evaluate whether the approach achieves the transparency of fuzzy rules without sacrificing the performance of modern policy-gradient RL.
2 Related Work
Recently, there has been growing interest in scaling up fuzzy RL and integrating it with deep learning. Zander et al. trained Takagi–Sugeno–Kang (TSK) fuzzy systems with Deep Q-Learning (DQN), reporting CartPole-v1 performance on par with, or better than, ordinary DQNs yet exhibiting the training instabilities typical of off-policy methods [6]. This motivates exploring on-policy optimisation such as Proximal Policy Optimisation (PPO).
3 Methodology
To isolate the effect of the optimization algorithm, we replicate the experimental setup of Zander et al., replacing their DQN learner with a PPO-based actor–critic loop. Four agents, each initialized with a distinct random seed, are trained on the CartPole-v1 environment.
As in the original work, the raw state vector is passed through a neural network comprising an input layer with 4 units, a hidden layer of 128 neurons with ReLU activation, and a second hidden layer of 127 ReLU-activated neurons. These 127 intermediate features are used to activate 16 Gaussian membership functions, whose firing strengths are used in a weight sum of the first-order TSK rule consequents to produce the final action logits.
Figure 1: Architecture of the controller
The value function, used to estimate state values for PPO’s critic, is modeled by a separate neural network. This network consists of two fully connected hidden layers with 64 and 32 units respectively, each followed by a Tanh activation.
Figure 2: Architecture of the value function
The actor and critic networks are optimized jointly using the PPO objective, which balances clipped policy updates, value regression, and entropy regularization to ensure stable learning. The full training loop for the PPO-ANFIS agent is summarized in Algorithm 1.
Algorithm 1 PPO-ANFIS training loop
1:Initialise ANFIS policy $\pi_{\theta}$ and value network $V_{\phi}$
2:for each iteration $i=1,\dots,N_{\text{updates}}$ do
3:     Collect $T$ timesteps of on-policy data $\mathcal{D}_{i}$ using $\pi_{\theta}$
4:     Compute returns $R_{t}$ and advantages $\hat{A}_{t}$ for all $(s_{t},a_{t})\in\mathcal{D}_{i}$
5:     for epoch $k=1$ to $K$ do
6:         Sample minibatch $\mathcal{B}\subset\mathcal{D}_{i}$
7:         $\mathcal{L}_{\text{clip}}=\mathbb{E}_{\mathcal{B}}\!\Big{[}\min\!\big{(}r_{t}(\theta)\hat{A}_{t},\;\operatorname{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\big{)}\Big{]}$
8:         $\mathcal{L}_{\text{VF}}=\frac{1}{2}(R_{t}-V_{\phi}(s_{t}))^{2}$(Value loss)
9:         $\mathcal{L}=-\mathcal{L}_{\text{clip}}+c_{v}\mathcal{L}_{\text{VF}}-c_{e}\,\mathbb{E}_{\mathcal{B}}[H[\pi_{\theta}(\cdot|s_{t})]]$
10:         Update $(\theta,\phi)$ via Adam; clip gradient-norm to $10$
Notation. $r_{t}(\theta)=\pi_{\theta}(a_{t}|s_{t})/\pi_{\theta_{\text{old}}}(a_{t}|s_{t})$; $c_{v}$ and $c_{e}$ are value and entropy weights; $\epsilon=0.2$ is the PPO clip parameter.
All experiments use the CartPole-v1 environment with the episode length capped at 500 steps. We use four runs that employ the following seeds $\{9,\allowbreak 42,\allowbreak 109,\allowbreak 131\}$; each run trains for $1{\times}10^{5}$ mini-batch updates. The hyperparameters are included in the Appendix Appendix.
4 Results
Figure 3: Evaluation performance of PPO-trained ANFIS agents on CartPole-v1. Mean return computed over 10 deterministic episodes, averaged across seeds.
Figure 4: Evaluation performance of ANFIS-DQN and vanilla DQN agents on CartPole-v1 from [6].
5 Discussion
Figure 3 shows the training performance of PPO-trained ANFIS agents across four random seeds in the CartPole-v1 environment while Figure 4 shows the equivlanet results in ANFIS-DQN and DQN signified by blue and orange respectively. All PPO-trained ANFIS agents demonstrate rapid convergence, with mean episode returns achieving the environment’s maximum score of 500 within the first 20,000 mini-batch iterations. This consistent upward trajectory reflects the stability of PPO’s clipped objective when applied to fuzzy policies.
Despite some fluctuations particularly with seed 42 around the 40,000–60,000 range, the agents eventually recover and maintain high performance throughout the remainder of training. Importantly, the final performance across all seeds to the optimal score, indicating that PPO not only handles the hybrid structure of the fuzzy policy well but also remains robust to initialization variability.
These results are very different from those observed in DQN-based ANFIS training from prior work [6], where the mean reward is unstable over the course of 100000 training iterations. Our findings suggest that PPO offers a more reliable optimization strategy for fuzzy controllers and may offer the necessary stability to develop robust and general solutions for high dimensional environments.
6 Future Work
In future work, we aim to expand and test this framework to more complicated environments like LunarLander-v3-Continuous and Hopper-v1. We also plan to explore integration of interpretability tools like SHAP or LIME to attribute actions to specific fuzzy rules, guiding rule pruning and the discovery of the optimal rule count.
{credits}
6.0.1 Acknowledgements
The authors extend their sincere gratitude to the members of the AI Bio Lab at the University of Cincinnati for their invaluable discussions and collaborative efforts that facilitated the realization of this work. The authors especially thank Bharadwaj ’Ben’ Dogga for his thoughtful review of early drafts.
[…]
The authors have no competing interests to disclose.
Appendix
Parameter Value
Discount factor $\gamma$ $0.99$
Learning rate $1{\times}10^{-5}$
PPO clip $\epsilon$ $0.2$
Entropy coefficient $c_{e}$ $0.02$
Mini-batch size $64$
Roll-out horizon $T$ $2048$ steps
Gradient-norm clip $10$
[…]
Membership widths $\sigma_{i}$ $0.25+0.5\,\mathcal{U}(0,1)$
Rule consequents $w_{i}$ $2\,\mathcal{N}(0,1)$

### v2

@@
 Abstract
-We propose a reinforcement learning (RL) approach for training neuro-fuzzy controllers using Proximal Policy Optimization (PPO).
-Building on prior work that applied Deep Q-Learning to Adaptive Neuro-Fuzzy Inference Systems (ANFIS), our method replaces the off-policy value-based framework with a stable on-policy actor–critic loop.
-We evaluate this approach in the CartPole-v1 environment using multiple random seeds and compare its learning performance against ANFIS-Deep Q-Network(DQN) baselines.
-It was found that PPO-trained fuzzy agents achieved a mean return of $500\pm 0$ on CartPole-v1 after 20000 updates, showcasing less variance than prior DQN-based methods during training and overall faster convergence.
-These findings suggest that PPO offers a promising pathway for training explainable neuro-fuzzy controllers in reinforcement learning tasks.
+We present a reinforcement learning method for training neuro-fuzzy controllers using Proximal Policy Optimization (PPO).
+Unlike prior approaches that used Deep Q-Networks (DQN) with Adaptive Neuro-Fuzzy Inference Systems (ANFIS), our PPO-based framework leverages a stable on-policy actor-critic setup.
+Evaluated on the CartPole-v1 environment across multiple seeds, PPO-trained fuzzy agents consistently achieved the maximum return of $500$ with zero variance after $20{,}000$ updates, outperforming ANFIS-DQN baselines in both stability and convergence speed.
+This highlights PPO’s potential for training explainable neuro-fuzzy agents in reinforcement learning tasks.
 Keywords:
@@
 In contrast, fuzzy inference systems offer transparency while providing a robust solution.
-These systems fall into two families: Mamdani and Takagi–Sugeno–Kang (TSK).
+These systems fall into two families: Mamdani and Takagi-Sugeno-Kang (TSK).
 Mamdani systems rely on linguistic IF–THEN rules with fuzzy outputs and subsequent defuzzification, making them highly interpretable but less amenable to gradient-based tuning [2].
-TSK models instead express rule consequents as algebraic functions of the inputs, producing smoother numeric outputs and enabling more robust numerical optimization [3].
+TSK models instead express rule consequents as linear functions of the inputs, producing smoother numeric outputs and enabling more robust numerical optimization [3].
 Yet both architectures still lack systematic training pipelines.
@@
 trained Takagi–Sugeno–Kang (TSK) fuzzy systems with Deep Q-Learning (DQN), reporting CartPole-v1 performance on par with, or better than, ordinary DQNs yet exhibiting the training instabilities typical of off-policy methods [6].
-This motivates exploring on-policy optimisation such as Proximal Policy Optimisation (PPO).
-3 Methodology
+This motivates exploring on-policy optimization such as PPO.
+3 Methodologies
 To isolate the effect of the optimization algorithm, we replicate the experimental setup of Zander et al., replacing their DQN learner with a PPO-based actor–critic loop.
@@
 As in the original work, the raw state vector is passed through a neural network comprising an input layer with 4 units, a hidden layer of 128 neurons with ReLU activation, and a second hidden layer of 127 ReLU-activated neurons.
-These 127 intermediate features are used to activate 16 Gaussian membership functions, whose firing strengths are used in a weight sum of the first-order TSK rule consequents to produce the final action logits.
+These 127 intermediate features are used to activate 16 Gaussian membership functions, whose firing strengths are used in a weight sum of the first-order TSK rule consequents to produce the final action.
 Figure 1: Architecture of the controller
@@
 We use four runs that employ the following seeds $\{9,\allowbreak 42,\allowbreak 109,\allowbreak 131\}$; each run trains for $1{\times}10^{5}$ mini-batch updates.
-The hyperparameters are included in the Appendix Appendix.
+The hyperparameters are included in the Appendix.
 4 Results
-Figure 3: Evaluation performance of PPO-trained ANFIS agents on CartPole-v1. Mean return computed over 10 deterministic episodes, averaged across seeds.
-Figure 4: Evaluation performance of ANFIS-DQN and vanilla DQN agents on CartPole-v1 from [6].
+(a) PPO-trained ANFIS agents on CartPole-v1. Mean return over 10 deterministic episodes, averaged across seeds.
+(b) Same experiment with gradient-norm clipping at 0.5.
+Figure 3: Evaluation curves for PPO-trained ANFIS controllers.
+Figure 4: ANFIS-DQN and vanilla DQN agents on CartPole-v1 (Referenced from [6]).
 5 Discussion
-Figure 3 shows the training performance of PPO-trained ANFIS agents across four random seeds in the CartPole-v1 environment while Figure 4 shows the equivlanet results in ANFIS-DQN and DQN signified by blue and orange respectively.
-All PPO-trained ANFIS agents demonstrate rapid convergence, with mean episode returns achieving the environment’s maximum score of 500 within the first 20,000 mini-batch iterations.
-This consistent upward trajectory reflects the stability of PPO’s clipped objective when applied to fuzzy policies.
-Despite some fluctuations particularly with seed 42 around the 40,000–60,000 range, the agents eventually recover and maintain high performance throughout the remainder of training.
-Importantly, the final performance across all seeds to the optimal score, indicating that PPO not only handles the hybrid structure of the fuzzy policy well but also remains robust to initialization variability.
-These results are very different from those observed in DQN-based ANFIS training from prior work [6], where the mean reward is unstable over the course of 100000 training iterations.
-Our findings suggest that PPO offers a more reliable optimization strategy for fuzzy controllers and may offer the necessary stability to develop robust and general solutions for high dimensional environments.
+Figure 3(a) shows the training performance of PPO-trained ANFIS agents in the CartPole-v1 environment across four random seeds, while Figure 4 presents the results for ANFIS-DQN and standard DQN.
+All PPO-based agents converge rapidly, reaching the maximum return of 500 within 20,000–40,000 mini-batch iterations.
+Some fluctuations remain (seeds 9 and 131 around 50,000 iterations), partly due to hyperparameter choices.
+With more optimal settings such as gradient-norm clipping at 0.5, training stability improves as shown in Figure 3(b).
+By 100,000 iterations, all PPO agents consistently achieve the maximum return, demonstrating robustness to initialization and effective handling of the fuzzy policy structure.
+In contrast, prior DQN-based ANFIS training [6] showed persistent instability.
+These results highlight PPO’s suitability for training fuzzy controllers and its potential for scalable, stable learning in higher-dimensional tasks.
 6 Future Work
@@
 The authors extend their sincere gratitude to the members of the AI Bio Lab at the University of Cincinnati for their invaluable discussions and collaborative efforts that facilitated the realization of this work.
-The authors especially thank Bharadwaj ’Ben’ Dogga for his thoughtful review of early drafts.
+The authors especially thank Bharadwaj ‘Ben’ Dogga for his thoughtful review of early drafts.
 […]
 The authors have no competing interests to disclose.
-Appendix
-Parameter Value
-Discount factor $\gamma$ $0.99$
-Learning rate $1{\times}10^{-5}$
-PPO clip $\epsilon$ $0.2$
-Entropy coefficient $c_{e}$ $0.02$
-Mini-batch size $64$
-Roll-out horizon $T$ $2048$ steps
-Gradient-norm clip $10$
-[…]
-Membership widths $\sigma_{i}$ $0.25+0.5\,\mathcal{U}(0,1)$
-Rule consequents $w_{i}$ $2\,\mathcal{N}(0,1)$